OpenStudy (anonymous):

three times x plus 4 times y is 1 more than 6 times x what is x

OpenStudy (anonymous):

x=54

OpenStudy (anonymous):

to clarify \[3x + 4y = 1 + 6x\] subtract both sides by 3x to get \[4y = 1 + 6x - 3x\] simplify to get \[4y=1+3x\] now subtract a 1 on both sides and solve for x